OpenStudy (happpily):

You have a choice two cell phone plans. Plan A: 39.99 per month for 600 minutes plus .40 for additional minute Plan B: $45 per month with unlimited minutes How many minutes must you talk in one month to make plan B the best cell phone plan for you?

OpenStudy (michele_laino):

hint: If I talk for \(N\) l minutes, in one month, then using plan A, I pay this: \[\Large 39.99 + \left( {0.4 \times N} \right)\] Plan B is the best plan if the subsequent condition holds: \[\Large 39.99 + \left( {0.4 \times N} \right) > 45\]

OpenStudy (michele_laino):

oops.. \(N\) represents the number of additional minutes

OpenStudy (happpily):

Is .4 suppose to be .40? Thank you so much for helping me btw !!!

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

yes! It is \($0.40\)

OpenStudy (happpily):

Is N 15?

OpenStudy (michele_laino):

I got \(N \simeq 13\)

OpenStudy (happpily):

Really? But why 13

OpenStudy (michele_laino):

here is why: \[\Large \frac{{45 - 39.99}}{{0.40}} = 12.525 \simeq 13\]

OpenStudy (michele_laino):

please try to solve this equation for \(N\): \[\Large 39.99 + \left( {0.4 \times N} \right) = 45\]

OpenStudy (happpily):

Sorry I'm so confused?! What's the first step I have to solve?

OpenStudy (happpily):

I'm Trying to write everything down in order

OpenStudy (michele_laino):

as first step, you have to solve this equation: \[\Large 39.99 + \left( {0.4 \times N} \right) = 45\] please find \(N\)

OpenStudy (happpily):

To find N I have to subtract 45 with 39 right?

OpenStudy (michele_laino):

that's right! \(45-39.99=...?\)

OpenStudy (happpily):

5.01!! :)

OpenStudy (michele_laino):

now, please divide by \(0.40\)

OpenStudy (happpily):

12.525

OpenStudy (michele_laino):

correct! it means if I talk for at least 13 additional minutes the cost of Plan A is equal to the cost of Plan B Whereas if I talk for additional 15 minutes or 14 minutes, then cost of Plan A is Greater than the cost of Plan B

OpenStudy (happpily):

Wait so 12.525 is minutes or cost.

OpenStudy (happpily):

Do we round ? How is 12.525 = 13 :0